OpenStudy (anonymous):

On the Apollo 14 mission to the moon, astronaut Alan Shepard hit a golf ball with a 6 iron. The acceleration due to gravity on the moon is 1/6 of its value on earth. Suppose he hits the ball with a speed of 27m/s at an angle 50 above the horizontal. a) How much farther did the ball travel on the moon than it would have on earth? b) For how much more time was the ball in flight?

OpenStudy (anonymous):

take t=3.5s, the velocity of x=dx/dt=3/2t^2-4t=4.375 velocity of y=dy/dt=t-2=1.5 the net velocity is obvious then speed=sqrt(Vx^2+Vy^2)=sqrt(4.375^2-1.5^2)=4.11 perhaps there's some mistake, check it. Anyway you should have understood

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

It's wrong, it says the answer should be in two significant figures

OpenStudy (anonymous):

it could be some calculative problem you can sort it out

OpenStudy (anonymous):

I got 5.875 still wrong thanks for trying

OpenStudy (anonymous):

4.625?

OpenStudy (anonymous):

21.39

OpenStudy (anonymous):

no it has to be two significant figures fore example: 45 , 67 not just one significant figure like 4 or 5

OpenStudy (anonymous):

got it1369/64

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

sorry,wait a minute

OpenStudy (anonymous):

I only got when t =0 the answer is 2 , but I don't know how to get t = 3.5 !!

OpenStudy (anonymous):

you don't use calculous?

OpenStudy (anonymous):

first derivative,than The Pythagorean theorem. that cannot be wrong

OpenStudy (anonymous):

I know that that's how I got when t =0 , I did the same exact way when t = 3.5 and I got 5.375 but wrong

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

my answer is 21.390625. I'm closely sure that it's correct

OpenStudy (anonymous):

\[x= (1/2t^{3} - 2t^{2})........ y= (1/2t^{2}-2t)....\] i just want to see the equations in a better format :D

OpenStudy (anonymous):

\[dx/dt= (-3/2t^{4} - 4t^{2}) ...... dy/dt = (-1/t^{3} - 2)\]

OpenStudy (anonymous):

now put t = 3.5

OpenStudy (anonymous):

that's wrong

OpenStudy (anonymous):

what ?? whats wrong in that?

OpenStudy (anonymous):

U should get the derivative not the integration of the equation

OpenStudy (anonymous):

omg.. !! please.. you need to work more on your math :-/